OpenStudy (anonymous):

Can someone show me examples of Associative Property of Addition Commutative Property of Addition Inverse Property of Addition

OpenStudy (anonymous):

www.youtube.com/watch?v=pB_zN9puUek Commutative property: When two numbers are added, the sum is the same regardless of the order of the addends. For example 4 + 2 = 2 + 4 Associative Property: When three or more numbers are added, the sum is the same regardless of the grouping of the addends. For example (2 + 3) + 4 = 2 + (3 + 4) Additive Identity Property: The sum of any number and zero is the original number. For example 5 + 0 = 5. Distributive property: The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example 4 * (6 + 3) = 4*6 + 4*3